# Pseudo-Stereo Inputs: A Solution to the Occlusion Challenge in Self-Supervised Stereo Matching

### v1

@@ III-A3 Method for Generating Pseudo-Images
While the most common warping operation in stereo matching involves using $D^{L}$, which is aligned with $I^{L}$, to warp $I^{R}$ into an image that aligns with $I^{L}$, this approach cannot be used for generating pseudo-images. The reason is that we do not have a disparity map aligned with the pseudo-image prior to its generation. Consequently, our approach employs a pseudo-image generation method that more closely resembles a simplified rendering process as illustrated in Figure 4. For $I^{R}$, each pixel can be relocated backward using $D^{R}$ to generate $\widetilde{I}^{R}$. If multiple pixels map to the same location in $\widetilde{I}^{R}$, only the pixel with the maximum disparity is kept. This corresponds to the occlusion effect during rendering. As shown in Figure 5, this process yields a pseudo-image containing missing pixels. To make it as similar to the real image as possible, it is necessary to pad these missing pixels. The padding method and how closely the resulting image resembles a real one directly influence the severity of the overfitting problem. This issue is discussed in detail in Section III-B.
III-B Feedback Conflicts and Overfitting Problems
Simply applying the pseudo-stereo inputs strategy with no further modifications already results in network performance that far surpasses current direct self-supervised methods, which can be evident from the ablation experiments in Section IV-B. Nevertheless, there remains a considerable gap when compared to dataset generation-based methods [13]. Our further experiments observed that feedback conflicts and overfitting problems are the reasons hindering further performance improvements. Specifically, as training progresses, the prediction in occluded regions becomes unstable, and the matching accuracy shows a declining trend towards the later stages of training.
III-B1 Feedback Conflicts
@@
This phenomenon can be readily explained. Despite our strategy in self-supervised feedback ensuring a 50% probability of obtaining correct information on the occluded side, there remains an equal 50% probability of receiving incorrect information. Although the information is always conflicting, it is expected that the network will attempt to minimize the loss of both correct and incorrect information simultaneously, resulting in a compromised outcome in the occluded region. As the gradient of photometric loss is negatively correlated with image similarity, correct disparity estimations receive lower gradients due to high similarity in later training stages. Conversely, incorrect estimations obtain higher gradients due to low similarity, thus dominating the optimization process. This phenomenon contributes to the instability observed in the occluded region.
@@
Occlusion estimation and detaching the propagation of loss from occluded pixels offers a promising solution to this challenge. While consistency check [24] and backward warping [25] are prevalent methods for occlusion estimation, they necessitate estimating the disparity map from the other viewpoint. This dependency makes the unreliability of the other view’s disparity prediction severely detrimental to the accuracy of occlusion estimation. We implement a solution based on the rendering method described in Section III-A3 for direct occlusion estimation. Specifically, we identify invalid pixels during the rendering process – those falling outside the image boundaries or occluded by other pixels – and classify them as occlusion pixels. This method leverages the inherent properties of the current view rather than relying on information from another view to determine pixel validity. Consequently, it effectively minimizes the risk of misclassifying non-occlusion pixels and prevents erroneous feedback from propagating. After mitigating erroneous feedback stemming from occluded pixels using this approach, we observe a notable improvement in the network’s performance. Detailed results can be found in the ablation study presented in Section IV-B. Subsequently, the photometric loss $L_{p}$ can be expressed as:
@@
where $\text{where }pe(I^{A},I^{B})$ $=\frac{\alpha}{2}(1-SSIM(I^{A},I^{B}))$ (2)
@@
Here, $p$ and $1-p$ represent the probabilities of the respective items, both being constant at $0.5$. $\widetilde{O}_{ij}$ denotes the occlusion estimation value, where it equals 1 for non-occluded pixels and 0 for occluded ones. $\alpha=0.85$ as in [12].
III-B2 Overfitting Problems
Incorporating occlusion estimation into the pseudo-stereo inputs strategy further enhances performance. Nevertheless, extended long-term training stability tests have revealed the overfitting problem. Further analysis of the experimental images reveals that this issue stems from the network’s ability to distinguish between real and pseudo images in the later stages of training. Specifically, in current strategies, the use of real or pseudo image pairs directly determines the location of occlusions. This enables the network to classify inputs by detecting whether they are pseudo images, thereby converging different results for each input type. This renders the network no longer “consistently have a probability to capture information from both sides of the occluders” as expected.
@@ III-B2 Overfitting Problems
These observations suggest that the generated pseudo-images contain sufficient features for the network to distinguish them. Based on this inference, we initially attempted to generate more indistinguishable images to prevent the network from classifying the inputs. When generating images using the rendering method described in Section III-A3, notable features include small-scale voids at occluded areas and large-scale missing areas along the right edge, as shown in Figure 5. For small-scale voids, we fill them using the average value of adjacent non-void pixels. For the missing parts at the right edge, we adopt a wider generation strategy to produce images that are wider than the inputs. This ensures that the resulting cropped images no longer suffer from large-scale missing regions at their edges. It is evident that the generated images now possess a higher degree of realism. Under this strategy, experiments demonstrate an improvement in overfitting issues; however, they have not been completely eradicated. This implies that the network is still able to extract subtle features for differentiation.
We further implement the fully pseudo-stereo inputs strategy. This strategy no longer relies on real image pairs but instead uses fully pseudo-stereo images as input. To be specific, the network’s inputs are changed to image pairs $I^{L}$,$\widetilde{I}^{L}$ and $I^{R}$, $\widetilde{I}^{R}$. At this point, since all the inputs are pseudo-stereo inputs, the network naturally cannot differentiate them further. As shown by the experiments in Section IV-B, the network achieves stable convergence to high performance and avoids overfitting under this strategy. At this point, (1) becomes,
@@
$\lambda$ is the smoothing term coefficient, starting from 0.001 and gradually increasing to the set value of 0.5 over 10k iterations.
TABLE I: Comparison with existing methods on the KITTI online benchmark, including representative supervised methods (S), self-supervised methods based on dataset generation (D), and direct self-supervised methods based on photometric consistency (P). The best results in each category are highlighted in bold.
@@
S Supervised methods - Training with ground truth
DispNetC [2] 4.11 4.65 0.9 1.0 4.32 4.41 4.34 4.05
PSMNet [3] 1.49 1.89 0.5 0.6 1.86 4.62 2.32 2.14
LoS [4] 1.10 1.38 0.4 0.4 1.42 2.81 1.65 1.52
D Self-supervised methods - Based on dataset generation
MfS-PSMNet [14] - [4.43] - [0.95] - - [4.75] [4.57]
NS-PSMNet [17] - - - - - - [5.05] [4.80]
Rversing-PSMNet [13] - - - - 3.13 8.7 4.06 3.86
P Direct self-supervised methods - Based on photometric consistency
@@ D Self-supervised methods - Based on dataset generation
Pseudo-Stereo (ours) 3.46 4.08 0.8 0.9 3.11 12.52 4.68 4.40
IV Experiments
In this section, we present experiments conducted with the methods proposed in this paper. Section IV-A compares our approach with existing supervised methods, data generation-based self-supervised methods and direct self-supervised methods. The ablation study is conducted in Section IV-B to demonstrate the contributions of each component described previously.
Dataset. In order to contrast with prior research [6, 8, 11, 22, 26], our experiments utilize the same datasets: KITTI 2012 [27] and KITTI 2015 [28]. All training processes are conducted entirely using the self-supervised method without any ground truth, relying solely on image pairs as both inputs and feedback. Each dataset mentioned above consists of a training set and a testing set. The training set provides ground truth data, while results on the test set must be submitted to their website for evaluation. Given that self-supervised training only requires image pairs, we utilize all the image pairs in each dataset for training. The training set, which provides ground truth, is used for validation and ablation studies. Test set results are submitted to their website for evaluation and comparison with existing methods.
@@ IV Experiments
Training Details. We maintain consistent parameters across different datasets to demonstrate the robustness of our method. All training is conducted using a single RTX 4090 GPU in PyTorch [29]. We employ the Adam optimizer [30] with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and the cosine annealing schedule. Following the settings in PSMNet [3], we crop images to $512\times 256$ during training, with a batch size of 8. Data augmentation is applied to the input images but not to the feedback images. This augmentation consists of random cropping, brightness enhancement, and contrast adjustment. Convergence to optimal performance occurs after approximately $35k$ iterations, with a total training time of about 7 hours. The training process starts from scratch without any pre-training or post-processing involved.
IV-A Comparisons with State-of-the-art
@@
In this section, our method is compared with other existing self-supervised methods [6, 8, 11, 22, 26, 14, 17, 13] and some representative supervised methods [2, 3, 4]. Given that all mentioned datasets offer clear online benchmarks, we first present the online benchmark results in Table I. The evaluation metrics are identical to those in the online benchmarks. Specifically, “noc” and “all” denote non-occluded pixels and all pixels respectively. “n-noc/all” refers to the percentage of pixels exceeding a difference of “n” from ground truth. “D1” indicates the proportion of pixels where the end-point error is less than 3px or 5%. “EPE” represents the average difference between estimated values and ground truth. Since some methods lack online test results, we include their reported training set results from the paper for reference only. Generally, these metrics tend to perform worse on the test set compared to the training set.
The results in Table I demonstrate that our method significantly outperforms other direct self-supervised methods across the board. When compared to dataset generation-based methods, we only slightly lag behind Rversing-PSMNet [13], which integrates complex operations such as traditional methods, deep completion networks, and label distillation. In addition, it is shown that our results have partially outperformed the earlier representative supervised method DispNetC [2].
Figure 6: Qualitative comparisons. Representative methods from different categories are selected, with all disparity maps sourced from the KITTI 2015 online benchmark. (S: supervised method; D: dataset generation-based method; P: direct self-supervised method based on photometric loss.)
In Figure 6, we select representative methods from different categories for qualitative comparison with our approach. Specifically, this includes the representative of supervised methods (S), Los [4]; the representative of dataset generation-based methods (D), Reversing [13]; and the representative of direct self-supervised learning methods based on photometric loss (P), PASMNet [8]. Among these methods, PASMNet, which belongs to the same category as ours, is notably impacted by occlusion problems and generates extensive erroneous predictions. In contrast, our method maintains accurate predictions in the same occluded regions. This suggests that the long-standing occlusion problem, which has been a hindrance in such methods, is addressed in our approach.
Further comparisons with the other two methods highlight our significant competitiveness even when compared to them. Furthermore, detailed comparisons reveal additional advantages of our method. Specifically, different training methods lead to distinct limitations and biases in their outputs. In Figure 6, rectangular boxes highlight the locations where errors occur in the comparative methods. Within the white boxes, it is shown that the supervised approach [4], heavily dependent on label data, outputs significantly broader edges because of erroneous ground truth. In the purple boxes, it is demonstrated that [13] exhibits excessive smoothing in ambiguous regions due to its reliance on labels generated by a monocular network. In contrast, our method does not rely on label data or data from other pre-processing methods. The direct self-supervised training based solely on images enables it to produce results closer to reality in the predicted disparity map. To further demonstrate this phenomenon, we select a typical scenario with problematic ground truth in Figure 1. In this scenario, supervised learning converges to incorrect results by strictly following the guidance of the ground truth. The previous self-supervised method fails to produce accurate results due to occlusion issues. Conversely, our self-supervised approach is unaffected by labeled data and also resolves occlusion problems, yielding more realistic predictions. This undoubtedly highlights the additional potential of our direct self-supervised method. Inadequately, as shown in Figure 1 (B) and (C), areas with highly indistinct colors and interference remain challenging for self-supervised approaches that rely on photometric consistency. The solution to this challenge could be crucial for advancing the framework introduced in this work.
IV-B Ablation Study
@@
In this section, the roles of the various components proposed in Section III within the overall framework are studied and discussed. The components involved include the pseudo-stereo inputs strategy, the more accurate pseudo-image generation method, the occlusion estimation, and the fully pseudo-stereo inputs strategy. The experimental results under different combinations of components are shown in Table II. In this context, “PS” denotes the use of the pseudo-stereo inputs strategy, “Wider” indicates wider input for more accurate image generation, “Occ” represents occlusion estimation, and “FPS” signifies the adoption of the fully pseudo-stereo inputs strategy.
TABLE II: Ablation study of components in the proposed method.
Method PS Wider Occ FPS D1-all (%) EPE-all Remarks
(a) base 5.67 1.17
(b) base + occ 5.73 1.24 Over-smoothing in occlusion areas
(c) Only pseudo-stere 4.99 1.07 Overfitting
(d) PS + fully pseudo strategy 5.17 1.11 /
(e) PS + Occlusion detection 4.08 0.99 Overfitting
(f) Only no wider generation 4.31 1.04 /
(g) Only no fully pseudo 4.11 0.99 Slight overfitting
(h) Pseudo-Stereo (full) 4.06 1.01 /
In the base method (a), common photometric loss is used, while other components such as smoothness loss, data augmentation methods, and training parameters are identical to those in the complete method. Obviously, the complete method (h) shows a significant improvement compared to the base method (a), in which none of the proposed strategies are applied. The contrast between (a) and (c) shows that merely implementing the pseudo-stereo inputs strategy leads to a substantial improvement in the base method. The comparisons between (c) and (e), as well as between (d) and (f), demonstrate that the introduction of occlusion estimation is crucial for performance improvement, regardless of whether the fully pseudo-stereo inputs strategy is adopted. The results from (c), (e), and (g) collectively indicate that overfitting is difficult to address without the fully pseudo-stereo inputs strategy. It is noteworthy that the comparison between (e) and (g) indicates more realistic and indistinguishable generated images can alleviate overfitting problems. In Figure 7, we plot the training process curves to visually illustrate this phenomenon. It is evident that the overfitting problem disappears when adopting the fully pseudo-stereo inputs strategy in (h). Finally, by comparing (e), (f), and (h), it can be inferred that using the fully pseudo-stereo inputs strategy will reduce accuracy to some extent when generated images exhibit noticeable defects. Conversely, when using a wider generation to produce more accurate images, this issue is effectively resolved. In general, within the overall framework, each component plays a distinct role. Through the combination of all components, we obtained highly accurate and stable training outcomes.
Figure 7: Illustration of overfitting problems. Following the implementation of the fully pseudo-stereo inputs strategy (h), the upward trend of D1-all is no longer observed.
V Conclusion

### v2

@@ 3.1.2 Method for Generating Pseudo-Images
While numerous methods for novel view synthesis exist [14, 24, 13], they typically require additional data to train the models themselves. Incorporating such methods would introduce extraneous data and priors, which contradicts the fundamental principle of self-supervised learning. Therefore, our approach must exclusively utilize information available within the original training pipeline. In the standard stereo estimation pipeline, the commonly used warping operation is a form of image synthesis; this involves using $D^{L}$, which is aligned with $I^{L}$, to warp $I^{R}$ into an novel image that aligns with $I^{L}$. However, this approach cannot be used for generating pseudo-images. The reason is that we do not have a disparity map aligned with the pseudo-image prior to its generation. Consequently, our approach employs a pseudo-image generation method that more closely resembles a simplified rendering process as illustrated in Fig. 4. For $I^{R}$, each pixel can be relocated backward using $D^{R}$ to generate $\widetilde{I}^{R}$. If multiple pixels map to the same location in $\widetilde{I}^{R}$, only the pixel with the maximum disparity is kept. This corresponds to the occlusion effect during rendering. As shown in Fig. 5, this process yields a pseudo-image containing missing pixels. To make it as similar to the real image as possible, it is necessary to pad these missing pixels. The padding method and how closely the resulting image resembles a real one directly influence the severity of the overfitting problem. This issue is discussed in detail in Sec. 3.2.
3.1.3 Discussion on Using Pseudo Images as Inputs
As described earlier, our method does not use a previously trained network to generate novel views. Instead, the generation of pseudo-images relies on the network that is currently undergoing training. In other words, images generated in the early stages of training, which severely lack realism, are also used as inputs during the training process. This raises concerns about the network’s convergence. However, in fact, forcing the network to exploit monocular cues during stereo training is a common and effective practice. As an example, in the supervised stereo matching method [27], random patches are used to occlude the right view to accomplish this objective.
The cornerstone of our method’s stable convergence is that the left input and the feedback image are always real and reliable. The effective feedback compels the network to leverage monocular cues in the initial stages of training to enhance prediction accuracy. As accuracy increases, the precision of pseudo-image generation also improves, leading to better binocular information for the network. This results in a stable and progressively enhancing iterative process. The experimental results in Section 4.3 also support this point. In these experiments, even when no real images were used for the right input (i.e., using $I^{L}$, $\widetilde{I}^{L}$ or $I^{R}$, $\widetilde{I}^{R}$ as network inputs), the network still converges successfully.
Figure 4: Schematic diagram of the pseudo-images generation using the rendering method, where Backbone is the model currently under training rather than a previously well-trained one. The subscript “Flip” denotes horizontal flipping.
3.2 Feedback Conflicts and Overfitting Problems
Simply applying the pseudo-stereo inputs strategy with no further modifications already enables a significant performance improvement over the standard paradigm, which can be evident from the ablation study in Sec. 4.3. However, the feedback conflicts and overfitting issues hinder further performance enhancements and prevent stable convergence.
3.2.1 Feedback Conflicts
@@
The feedback conflicts issue manifests as unstable disparity estimates in occluded regions. This phenomenon can be readily explained. Despite our strategy in self-supervised feedback ensuring a 50% probability of obtaining correct information on the occluded side, there remains an equal 50% probability of receiving incorrect information. Under the influence of this conflicting feedback, the network will attempt to minimize the loss of both correct and incorrect information simultaneously, finally resulting in a compromised outcome in the occluded region. More importantly, the gradient of the photometric loss is negatively correlated with image similarity. As training progresses and overall image similarity increases, the global average gradient diminishes to a low value. However, occluded regions, which lack correct correspondences, produce a large color discrepancy when warped even with the correct disparity.
@@
Occlusion estimation and detaching the propagation of loss from occluded pixels offers a promising solution to this challenge. While consistency check [9] is an established method for occlusion estimation, it suffers from two notable drawbacks. First, it incurs additional computational cost by requiring the inference of a disparity map for the opposing view. Second, it is prone to errors where incorrect disparities coincidentally satisfy the consistency check, a problem particularly prevalent at object boundaries. We implement a solution that leverages the inherent properties of the current view rather than relying on information from another view. Specifically, we identify invalid pixels during the rendering process—those falling outside the image boundaries or occluded by other pixels—and classify them as occlusion pixels. This method effectively minimizes the risk of misclassifying pixels and prevents erroneous feedback from propagating. After mitigating erroneous feedback stemming from occluded pixels using this approach, we observe a notable improvement in the network’s performance. Detailed results can be found in the ablation study presented in Sec. 4.3. Subsequently, the photometric loss $L_{p}$ can be expressed as:
@@
$\text{where }pe(I^{A},I^{B})$ $=\frac{\alpha}{2}(1-SSIM(I^{A},I^{B}))$ (2)
@@
Here, $p$ and 1p1-p represent the probabilities of the respective items, both being constant at $0.5$. $\widetilde{O}_{ij}$ denotes the occlusion estimation value, where it equals 1 for non-occluded pixels and 0 for occluded ones. $\alpha=0.85$ as in [7].
3.2.2 Overfitting Problems
Incorporating occlusion estimation into the pseudo-stereo inputs strategy further enhances performance. Nevertheless, extended long-term training stability tests have revealed the overfitting problem, manifested as a gradual decline in accuracy after reaching a peak. Critically, self-supervised training, by its nature, operates without labeled data and thus lacks a validation set. Consequently, early stopping cannot be relied upon to prevent overfitting; instead, the network must be guaranteed to converge stably.
We identify the root cause of this overfitting as the network’s ability to distinguish between real and pseudo images in the later stages of training. Specifically, in current strategies, the use of real or pseudo image pairs directly determines the location of occlusions. This enables the network to classify inputs by detecting whether they are pseudo images, thereby converging different results for each input type. This renders the network no longer “consistently have a probability to capture information from both sides of the occluders” as expected.
@@ 3.2.2 Overfitting Problems
These observations suggest that the generated pseudo-images contain sufficient features for the network to distinguish them. Based on this inference, we initially attempted to generate more indistinguishable images to prevent the network from classifying the inputs. When generating images using the rendering method described in Sec. 3.1.2, notable features include small-scale voids at occluded areas and large-scale missing areas along the right edge, as shown in Fig. 5. For small-scale voids, we fill them using the average value of adjacent non-void pixels. For the missing parts at the right edge, we adopt a wider generation strategy to produce images that are wider than the inputs. This ensures that the resulting cropped images no longer suffer from large-scale missing regions at their edges. It is evident that the generated images now possess a higher degree of realism. Under this strategy, experiments demonstrate an improvement in overfitting issues; however, they have not been completely eradicated. This implies that the network is still able to extract subtle features for differentiation.
We further implement the fully pseudo-stereo inputs strategy. This strategy no longer relies on real image pairs but instead uses fully pseudo-stereo images as input. To be specific, the network’s inputs are changed to image pairs $I^{L}$$\widetilde{I}^{L}$ and $I^{R}$, $\widetilde{I}^{R}$. At this point, since all the inputs are pseudo-stereo inputs, the network naturally cannot differentiate them further. As shown by the experiments in Sec. 4.3, the network achieves stable convergence to high performance and avoids overfitting under this strategy. At this point, (1) becomes,
@@
$\lambda$ is the smoothing term coefficient, starting from 0.001 and gradually increasing to 0.5 over 10k iterations.
Table 1: Comparison with existing methods on the KITTI online benchmark.
@@
Pseudo-Stereo-PSMNet (ours) 3.46 4.08 0.8 0.9 3.11 12.52 4.68 4.40
Pseudo-Stereo-MonSter (ours) 3.08 3.62 0.7 0.8 2.93 11.67 4.39 4.22
4 Experiments
In this section, we present experiments to validate that our method successfully resolves the occlusion problem. The logic of our experimental validation is as follows. First, in Sec. 4.1, we conduct a direct comparison against existing methods on the mainstream benchmark datasets they established. Then, we perform additional experiments across various datasets and backbone networks to verify the generalizability of our method in Sec. 4.2. Finally, Sec. 4.3 presents ablation studies to validate the contribution of each proposed component. We emphasize the visualization of occluded regions to confirm the resolution of the occlusion problem, while also showcasing the performance improvements achieved by solving it.
Datasets and Backbones. For evaluation, we use the KITTI 2012 [5] and KITTI 2015 [23] datasets, which consist of real-world driving scenes and are the established benchmarks for comparing with previous methods [16, 31, 35, 19, 3]. To validate the robustness of our method, we also use two additional datasets with vastly different styles: SceneFlow [20] and Spring [21], with the latter being used at half resolution owing to our limitations in computational resources. For the network architectures, we select three representative models with highly distinct designs—PSMNet [1], GMStereo [34], and MonSter [2]—to verify the general applicability of our method across different backbones.
@@ 4 Experiments
Training Details. We maintain consistent parameters across different datasets to demonstrate the robustness of our method. All training is conducted using two RTX 4090 GPUs in PyTorch [25]. We employ the Adam optimizer [15] with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and the cosine annealing schedule. Following the settings in [1], we crop images to $512\times 256$ during training, with a batch size of 8. Data augmentation is applied to the input images but not to the feedback images. This augmentation consists of random cropping, brightness enhancement, and contrast adjustment. All training processes start from scratch without any pre-training, ground-truth labels, or post-processing involved.
4.1 Comparisons with State-of-the-art
@@
We first present our submitted results on the KITTI online leaderboard in Tab. 1. The evaluation metrics are identical to those in the online benchmarks. Specifically, “noc” and “all” denote non-occluded pixels and all pixels respectively. “n-noc/all” refers to the percentage of pixels exceeding a difference of “n” from ground truth. “D1” indicates the proportion of pixels where the end-point error is more than 3px and 5%. “EPE” represents the average difference between estimated values and ground truth.
The results in Tab. 1 demonstrate that our method significantly outperforms other methods across the board. We present results using two distinct backbones: PSMNet [1] and MonSter [2]. The former is chosen to ensure that the performance gains are attributable to our method rather than to advancements in network architecture, thereby validating the credibility of our contribution. The latter validates the generalizability and applicability of our method to the latest architectures. These performance improvements are attributed to our successful resolution of the occlusion problem, which is visually demonstrated in Fig. 6. The figure presents representative occlusion scenarios alongside the disparity and error maps from different methods. It is clearly visible that other methods exhibit varying degrees of error on the left side of the occluder, which indicates that the occlusion problem remained unresolved in these prior works. In contrast, only our method produces results that are symmetric and accurate on both sides of the occluder. This validates that the primary objective of this work—to solve the occlusion challenge—has been successfully achieved.
Figure 6: Qualitative comparisons. All disparity and error maps sourced from the KITTI 2015 online benchmark. Only our method produces accurate results on both sides of occluders, demonstrating that the occlusion problem is resolved.
4.2 Evaluation on Various Backbones and Datasets
This section aims to demonstrate the robustness and effectiveness of our approach across various backbone architectures and datasets. As shown in Fig. 7, the occlusion problem is prevalent across various datasets, manifesting as erroneous disparity inference on the left side of foreground occluders (indicated by red pixels in the error maps). The proposed Pseudo-Stereo strategy demonstrates clear effectiveness in these occluded regions, yielding accurate results on both sides of the occluder. This qualitative improvement translates to quantitative gains, as shown in Tab. 2, where applying our method to different backbones on various datasets all leads to substantial accuracy enhancements attributable to solving the occlusion problem.
Figure 7: Qualitative results of our proposed Pseudo-Stereo method on the SceneFlow [20] and Spring [21] datasets, demonstrating its ability to resolve the occlusion challenge across various scenarios.
4.3 Ablation Study
@@
In this section, the roles of the various components proposed in Sec. 3 within the overall framework are studied and discussed. All evaluations are performed on the KITTI 2015 training set. The components involved include the pseudo-stereo inputs strategy, the refined pseudo-image generation method, the occlusion estimation, and the fully pseudo-stereo inputs strategy. The experimental results are shown in Tab. 3.
The results for settings (a-d) indicate that both the pseudo-stereo inputs and occlusion estimation are essential for performance enhancement, but they concurrently induce overfitting. Conversely, the results from (e-g), where overfitting is absent, prove that our fully pseudo-stereo input strategy is the key to resolving the overfitting problem. In Fig. 8, we plot the training curves to visually illustrate this phenomenon. The comparison between (f) and (g) reveals that when the generated pseudo-image suffers from significant artifacts, adopting the fully pseudo-stereo strategy can lead to a degradation in accuracy. This issue is effectively resolved when our refined pseudo-image generation strategy is used. Overall, each component plays an indispensable role within the entire framework. The combination of all components enables our model to achieve both high accuracy and stable training.
Table 2: Evaluation on various backbones and datasets.
Backbone w/o Pseudo-Stereo w/ Pseudo-Stereo
EPE (px) $\downarrow$ 3-px (%) $\downarrow$ EPE (px) $\downarrow$ 3-px (%) $\downarrow$
SceneFlow [20] Dataset
PSMNet [1] 5.41 16.85 3.10 9.80
GMStereo [34] 5.09 15.01 3.11 8.49
MonSter [2] 5.28 16.15 2.32 7.86
Spring [21] Dataset
PSMNet [1] 2.41 6.58 2.30 5.12
GMStereo [34] 2.33 6.31 2.18 4.78
MonSter [2] 2.37 6.56 1.77 4.65
Table 3: Ablation study of our proposed components. PS: Pseudo-stereo inputs strategy. Occ: Occlusion estimation and gradient detaching. RG: Refined pseudo-image generation. FPS: Fully pseudo-stereo inputs strategy.
Config. PS Occ RG FPS D1 (%) EPE Remarks
(a) 5.67 1.17 /
(b) 5.73 1.24 Over-smoothing in occlusions
(c) 4.99 1.07 Overfitting
(d) 4.08 0.99 Overfitting
(e) 5.17 1.11 /
(f) 4.31 1.04 /
(g) 4.06 1.01 /
Figure 8: Visualization of the overfitting problem. By adopting the fully pseudo-stereo input strategy in (f), the increasing D1 error trend observed in (c) is eliminated.
5 Conclusion